OpenStudy (anonymous):

rationalize the denominator of the expression. -5 over 1-√2

OpenStudy (anonymous):

\[\frac{-5}{1-sqrt[2]}\]

OpenStudy (anonymous):

5 rad 2 over -1

OpenStudy (saifoo.khan):

5+5root2

OpenStudy (anonymous):

\[\frac{-5}{1-\sqrt{2}}\times \frac{1+\sqrt{2}}{1+\sqrt{2}}\] is the method

OpenStudy (anonymous):

numerator is what the distributive law, denominator is \[1-2=-1\] so you get \[\frac{-5-5\sqrt{2}}{-1}=5+5\sqrt{2}\]

OpenStudy (anonymous):

what saifoo said

OpenStudy (saifoo.khan):

Yeah! :)